morganislame:

What is the equation, in point-slope form, for a line that goes through (8,−4) and has a slope of −56 ? y+4=−56(x−8) y+4=−56(x+8) y−4=−56(x+8) y−4=−56(x−8)

Mehek:

Point Slope Form: \(\bf y-y_1=m(x-x_1)\) Where m is the slope, and \(\bf{x_1}\) and \(\bf{y_1}\) are the ordered pair. Your ordered pair is (8, -4) so x = 8 y = -4 Plug them in: \(\bf y-y_1=m(x-x_1)\\ y - (-4) = -56(x-(8)\ \\ y + 4 = -56(x + 8)\) So it is Option B)

563blackghost:

@Mehek it could not be B. You are correct on the input of the y-value but the input of the x-value is incorrect. It would be \({(x+8)}\) if 8 were \({-8}\) but it is not. So this would mean that it would be placed as..\(\bf{(x-8)}\). So the true point-slope form is...\(\bf{y+4=-56(x-8)}\)

Mehek:

@563blackghost I did not notice the error, but ty for pointing it out

Vocaloid:

Closed due to answered.